OpenStudy (anonymous):

the length of a rectangle is twice the width. the area is 98yd^2. find the length and width

OpenStudy (anonymous):

Length*Width=98yd^2 Length:Width=2:1

OpenStudy (anonymous):

\[98yd \times 2/3\] is equal to the length

OpenStudy (anonymous):

and \[98yd \times 1/3\] is the width

OpenStudy (saifoo.khan):

Back! :D

OpenStudy (anonymous):

so length = 65.3333333... and width =32.6666666.... and if you want to double check 65/2=32.5 32*2=64

OpenStudy (anonymous):

hi saifoo!

OpenStudy (anonymous):

don;t forget to give me feedback :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

got ya joker, thanks!

OpenStudy (saifoo.khan):

the fun part is, i dont really remember that i helped you with e. LOL.

OpenStudy (anonymous):

you were in the middle of helping me when your computer conked lol

OpenStudy (saifoo.khan):

Ohhh! i guess i remember you now!!!